OpenStudy (anonymous):

(x-3)^2 Pleasee Help :)

hartnn (hartnn):

\[(a-b)^2=a^2+2ab+b^2\] put b=3

OpenStudy (anonymous):

huh? i dont understand

hartnn (hartnn):

that was the formula ,u just have to put b=3 like this: \[a^2+2*3*a+3^2\] now u just have to find 2*3 and 3^2=3*3....

OpenStudy (anonymous):

ohh okay how do you do that?

OpenStudy (anonymous):

nvm blond moment i know how to do that part

OpenStudy (anonymous):

a^2+6*a+9

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

i still have to combine right?

hartnn (hartnn):

thats your final answer a^2+6*a+9

OpenStudy (anonymous):

so you dont have to combine anhything?

OpenStudy (anonymous):

atleat the final asnwer should be a^2+15*9

hartnn (hartnn):

no u cannot different terms...6 has a with it,9 is constant.. u can combine like terms only.....like 2a+3a=5a and 10+20=30.....

OpenStudy (phi):

so you dont have to combine anhything? hartnn gave you a short cut. if you used FOIL on (x-3)(x-3) you would get 4 terms, 2 of which combine x^2-3x-3x+9 now combine the -3x -3x

OpenStudy (phi):

btw, that middle term should be negative.

OpenStudy (phi):

If you know how, I would do FOIL to get the answer. It is good practice.